Line segment RS=3x-3 and Line Segment ST=2x+6. What is the value of line segment RS if ST is twice the length of RS?
Answer by stanbon(75887)

Line segment RS=3x-3 and Line Segment ST=2x+6. What is the value of line segment RS if ST is twice the length of RS?
----------------------
Draw the picture of R.....S................T
----
Equation:
ST = 2(RS)
2x+6 = 2(3x-3)
2x+6 = 6x-6
4x = 12
x = 3
---
RS = 3x-3 = 3*3-3 = 6
===============================
Cheers,
Stan H.
=============